### v4

# Causal Panel Analysis under Parallel Trends:

Abstract
Two-way fixed effects (TWFE) models are widely used in political science to establish causality, but recent methodological discussions highlight their limitations under heterogeneous treatment effects (HTE) and violations of the parallel trends (PT) assumption. This growing literature has introduced new estimators and diagnostics, causing confusion among researchers about the reliability of existing results and best practices. To address these concerns, we replicated and reanalyzed 49 articles from leading journals using TWFE models for observational panel data with binary treatments. Using six HTE-robust estimators, diagnostic tests, and sensitivity analyses, we find: (i) HTE-robust estimators yield qualitatively similar but highly variable results; (ii) while a few studies show clear signs of PT violations, many lack evidence to support this assumption; and (iii) many studies are underpowered when accounting for HTE and potential PT violations. We emphasize the importance of strong research designs and rigorous validation of key identifying assumptions.
Keywords: panel data, two-way fixed effects, parallel trends, heterogeneous treatment effects, pretrend, event-study plot, difference-in-differences, robust confidence set
1.  Introduction
Over the past decade, political scientists have increasingly relied on panel data to draw causal conclusions (Xu, 2023). A favored method for such analyses is the two-way fixed effects (TWFE) model because of its ability to control for unobserved time-invariant confounders and common time trends. In our survey of 102 articles published from 2017 to 2023 in three top political science journals using observational panel data with dichotomous treatments, 64 studies (63%) assume a TWFE model with the following functional form or a close variant: [^1] The remaining 38 articles can be categorized into five groups: articles focusing on interaction effects (8 articles), articles using nonlinear links such as logit and Poisson (5 articles), articles employing instrumental variables or regression discontinuity designs (8 articles), and articles using other linear specifications, such as only one-way fixed effects or lagged dependent variables (17 articles). In some articles classified as using TWFE models, “unit” fixed effects are specified at the group level $g$, where multiple units $i$ are nested (e.g., county fixed effects when $i$ indexes cities), or time fixed effects are at a higher level $p$ (e.g., year fixed effects when $t$ indexes days). For simplicity, we use the notation $\alpha_{i}$ and $\xi_{t}$ rather than the more general $\alpha_{g}$ and $\xi_{p}$. For all $i,t$,
$Y_{i,t}=\delta^{TWFE}D_{i,t}+X_{i,t}{{}^{\prime}}\beta+\alpha_{i}+\xi_{t}+\varepsilon_{i,t},$ (1)
where $Y_{i,t}$ and $D_{i,t}$ are the outcome and treatment variables for unit $i$ at time $t$; $X_{i,t}$ is a vector of covariates; $\alpha_{i}$ and $\xi_{t}$ are unit and time fixed effect; and $\varepsilon_{i,t}$ is idiosyncratic errors. Researchers typically interpret $\delta^{TWFE}$ as the treatment effect and estimate the model using ordinary least squares. The resulting estimator for $\delta^{TWFE}$ is commonly known as the TWFE estimator. Moreover, researchers frequently conflate this model with a difference-in-differences (DID) design, and use the two terms interchangeably. [^2] We use the phrase “DID design” in reference to DID research design, which differs from the typical usage in the statistics literature that refers to treatment assignment mechanism.
Recent methodological discussions have raised concerns about the validity of TWFE models and the associated identifying assumptions, leaving many researchers in a quandary. First, existing findings based on the TWFE models may not hold given recent developments. Second, with the introduction of numerous new estimators and diagnostics, there is confusion about the current best practices. This paper seeks to bridge this gap by reviewing new estimation, inference, and diagnostics methods from the methodological literature and by reanalyzing published studies using both new estimators and the TWFE estimator. Based on the findings, we offer recommendations for researchers.
These criticisms of the use of TWFE models mainly come from two directions. First, the identification of TWFE models require the strict exogeneity assumption, which critics argue is stronger than many researchers realize and is often unrealistic in real-world settings (e.g., Imai and Kim, 2019). Strict exogeneity states that:
[…]
in which $\mathbf{D_{i}}=\{D_{i,1},D_{i,2},\cdots,D_{i,T}\}$ and $\mathbf{X_{i}}=\{X_{i,1},X_{i,2},\cdots,X_{i,T}\}.$ It means that once current treatment status, covariates, and fixed effects are accounted for, treatment status in any other periods has no partial effect on $Y_{i,t}$ (Wooldridge, 2010, p. 253). Under Model (1), strict exogeneity implies the parallel trends (PT) assumption:
(Parallel trends) $\mathbb{E}[Y_{i,t}(0)-Y_{i,s}(0)\mid D_{i,t}=1,D_{i,s}=0,X_{i,t}-X_{i,s}=x_{0}]$
=\ = $\mathbb{E}[Y_{j,t}(0)-Y_{j,s}(0)\mid D_{j,t}=0,D_{j,s}=0,X_{j,t}-X_{j,s}=x_{0}],\vspace{0.5em}$
in which $Y_{i,t}(0)=Y_{i,t}(d_{i,t}=0)$ represents the untreated potential outcome for unit $i$ at time $t$. It states that the change in untreated potential outcomes between any two periods is mean independent of the change in observed treatment status during those periods, once changes in covariate values are controlled for. Threats to PT, such as the presence of time-varying confounders and feedback from past outcomes to current treatment assignment, also invalidate strict exogeneity. Therefore, throughout the rest of the paper, we use the term “PT violations” to encompass violations of strict exogeneity. [^3] We discuss the relationship between strict exogeneity and PT, as well as other assumptions, under Model (1) in Section A.1.4 in the Supplementary Materials (SM). Note that the PT assumption invoked by many HTE-robust estimators does not depend on Model (1).
The second group of criticism concerns the consequences of heterogeneous treatment effects (HTE), i.e., $\delta^{TWFE}$ is not a constant (e.g., Goodman-Bacon, 2021; de Chaisemartin and D’Haultfœuille, 2020; Strezhnev, 2018; Sun and Abraham, 2021; Callaway and Sant’Anna, 2021; Athey and Imbens, 2022; Borusyak, Jaravel and Spiess, 2024). Researchers have shown that, under HTE, TWFE estimates in general do not converge to a convex combination of the individual treatment effects (ITE) for units under the treatment condition, even when the PT assumption is valid. The so-called “negative weighting” problem, as described in de Chaisemartin and D’Haultfœuille (2020), is an extremely alarming theoretical result because it implies that a TWFE estimand can be negative (positive) even when all ITE are positive (negative). To address this issue, researchers have proposed many new estimators that are “HTE-robust”—that is, estimators that converge to some convex combinations of ITE under their identifying assumptions.
This paper thus pursues two goals. First, we explain and compare six recent proposals to amend TWFE models, including the interaction weighted (IW) estimator (Sun and Abraham, 2021), stacked DID (Baker, Larcker and Wang, 2022), CS-DID (Callaway and Sant’Anna, 2021), DID multiple (de Chaisemartin and D’Haultfœuille, 2020), PanelMatch (Imai, Kim and Wang 2023, hereafter IKW, 2023), and the imputation method (Borusyak, Jaravel and Spiess 2024, hereafter BJS, 2024; Liu, Wang and Xu 2024, hereafter LWX, 2024). These estimators produce causally interpretable estimates under HTE and PT (or its variants). Second, we replicate and reanalyze 49 studies published in the American Political Science Review (APSR), American Journal of Political Science (AJPS), and The Journal of Politics (JOP) from 2017 to 2023 which rely on a TWFE model to draw causal conclusions. Our aim is to assess the consequences of using or not using HTE-robust estimators and shed light on the prevalence of PT violations in political science research.
Our reanalysis shows that there is considerable variation in the degree to which HTE-robust estimators agree with the TWFE estimator. In most studies, the estimators produce qualitatively similar estimates. A non-negligible minority, however, exhibit significant disagreements. In three cases, at least one of the HTE-robust estimators produces an estimate of the opposite sign as the TWFE estimate; for one of these cases, one estimate of the opposite sign is also statistically significant at 5%. There is also a more widespread problem of power: HTE-robust estimators tend to have larger measures of uncertainty, which, combined with even small fluctuations in point estimates, can weaken statistical confidence. This is especially relevant for results that originally teeter on the brink of significance.
The primary concern, however, is the validity of the PT assumption. While only a few studies show clear signs of PT violations, which likely lead to spurious findings, most studies lack the power to rule out that realistic PT violations could explain a non-zero estimated causal effect. In such cases, even mild PT violations prevent us from concluding that the original treatment effect is nonzero. This does not imply that these studies are incorrect; it simply means that the available data lack sufficient power to reject the null hypothesis of no effects when the PT assumption does not perfectly hold.
Overall, we find that about one-fourth of the articles in our sample meet our standard of being highly credible. In these articles, we can statistically distinguish the treatment effect from zero using an HTE-robust estimator, even allowing for mild PT violations benchmarked against placebo estimates using pre-treatment data. To be clear, we do not intend to criticize the authors of articles we have replicated, as many of the methods we employed were not available when their work was conducted. Our aim is to guide and improve future research.
Most importantly, we urge researchers to prioritize a strong research design and sufficient power in causal panel studies. Credible observational studies should feature a well-defined treatment-outcome pair, shock-induced variation in treatment assignment, and enough power to ensure results are not undermined by small perturbations of key identifying assumptions. Research design has often been overlooked in causal panel analyses, likely because researchers have become accustomed to accepting the strong assumptions behind TWFE models. Recent studies have emphasized the importance of (quasi-)randomness in treatment assignment for the robustness of DID findings (e.g., Roth and Sant’Anna, 2023). We offer recommendations on the choice of estimators and diagnostics in the final section.
This paper makes several contributions. First, we propose a typology of various estimators for causal panel analysis. Our typology is based on the settings in which an estimator can be used and how controls are chosen. We also provide a comprehensive comparison of these estimators and show how several proposals are actually equivalent in some circumstances. We hope this discussion helps researchers deepen their understanding of these estimators. Second, we adapt the robust confidence set (CS) approach for sensitivity analysis proposed by Rambachan and Roth (2023) to the setting of counterfactual estimators. We find it highly useful as it avoids the issue of conditional inference–where hypothesis testing conditional on passing a pretest (e.g., a pretrend test) often results in incorrect size (Roth, 2024). Third, our reanalysis both instills confidence in existing political science research conducted with TWFE models when done properly and raises warnings about potential risks, such as the failure of the PT assumption and a lack of power. Our work also contributes to the ongoing conversation on replication and reproducibility in political science (e.g., Eggers et al., 2015; Lall, 2016; Hainmueller, Mummolo and Xu, 2019; Lal et al., 2024).
Our work is closely related to Baker, Larcker and Wang (2022), who evaluate the credibility of a handful of studies with staggered adoption treatments in the finance and accounting literature. It differs in that: (i) we apply a broader range of estimators and diagnostic tests to a much larger number of empirical applications from a more diverse set of settings—notably, over half of these applications involve treatment reversal, where an additional no-carryover-effect assumption may be violated under TWFE models; (ii) our comprehensive review suggests that while weighting issues under HTE is important, the primary threats to causal inference with panel data, we believe, are PT violations and insufficient power. Our work is also related to McKenzie (2020), Roth et al. (2023), Xu (2023), and Arkhangelsky and Imbens (2024), who review and synthesize the recent methodological advancements of DID in the literature. What sets this paper apart is that the application of these innovations to data, which allows us to assess the relative importance of theoretical criticisms and evaluate the robustness of existing findings.
This research has a few limitations. First, we do not examine methods based on sequential ignorability, an alternative identification framework that assumes no unobserved confounding while allowing for dynamic treatment selection based on observables up to the current time period. Second, our analysis does not encompass studies that use continuous treatments, which is common in political science research. Finally, we note that the methodological literature is still developing, and thus our recommendations should be considered provisional as best practices evolve. Despite these limitations, which we aim to address in subsequent studies, we believe our replication and reanalysis exercise provides valuable insights into some of the widespread and significant challenges in using TWFE models for panel data analysis.
2.  TWFE and Its Pitfalls
In this section, we review the pitfalls of TWFE models identified in the literature. In the classic two-group and two-period case, the TWFE estimator $\hat{\delta}^{TWFE}$ is equivalent to the DID estimator, which consistently estimates the average treatment effect on the treated (ATT) under no anticipation and PT even with HTE. These results do not hold more generally in more complex settings with differential treatment adoption times (known as staggered adoption) or treatment reversal, as we will discuss below.
Our survey of the top journals reveals that the TWFE model under Model (1) is the most commonly adopted approach for estimating causal effects using panel data in political science. Fixed effects (FE) models began their rise to prominence in political science in the early 2000s, and criticism promptly followed. For example, in a debate with Green, Kim and Yoon (2001), Beck and Katz (2001) and King (2001) argue that linear FE models often lead to misleading findings because they throw away valuable information in data, ignore rich temporal dynamics, and are incapable of capturing complex time-varying heterogeneities. Moreover, because both the treatment and outcome variables are often serially correlated in a panel setting, researchers have cautioned against using standard error (SE) estimators suitable for cross-sectional data, such as Huber-White robust SEs (Bertrand, Duflo and Mullainathan, 2004). Scholars have also advocated for bootstrap procedures to better control Type I error rates when the number of clusters (units) is small (Cameron, Gelbach and Miller, 2008).
Recent criticisms.
In the past few years, a surge of studies has renewed investigation into the proprieties of the TWFE estimator and the assumptions it requires to achieve casual identification. One group of work urges researchers to better understand TWFE models (and their assumptions) from a design-based perspective, with a focus on restrictions on treatment assignment mechanisms. For example, Imai and Kim (2019) point out that the strict exogeneity assumption not only implies the well-known no time-varying confounder requirement, but it also forbids a “feedback” effect from past outcomes to treatment assignment. Blackwell and Glynn (2018) clarify that such an assumption is closely related to baseline randomization in which the treatment vector is generated prior to, or independent of, the realization of the outcome. Another body of work cautions researchers that the PT assumption, even in a simple $2\times 2$ DID setting, is sensitive to functional form. For example, Kahn-Lang and Lang (2020) emphasize the implicit functional form restrictions imposed by PT, encouraging researchers to justify it from a (quasi-)experimental perspective and address pre-treatment covariate imbalance. Roth and Sant’Anna (2023) point out that strong assumptions are needed for PT to be scale-independent, ensuring that a monotonic transformation of the outcome does not invalidate it. In practice, we find that many political science studies do not provide strong evidence to justify strict exogeneity or PT.
Figure 1: Toy Examples: TWFE Assumptions Satisfied vs. Violated
(a) Staggered DID: without Treatment Reversal
(b) Staggered DID: with Treatment Reversal
Note: The above figures show outcome trajectories of units in a staggered adoption setting (a) and in a general setting (b). Solid and hollow circles represent observed outcomes under the treatment and control conditions during the current period, respectively, while triangles represent counterfactual outcomes (in the absence of the treatment across all periods), $Y_{i,t}(\mathbf{d_{i}}=\mathbf{0})$. The data on the left panels in both (a) and (b) are generated by DGPs that satisfy TWFE assumptions while the data on the right are not. The divergence between hollow circles and triangles in the right panel of (b), both of which are under the control condition, is caused by carryover effects.
A second body of research explores the implications of HTE with binary treatments within TWFE models (e.g., Goodman-Bacon, 2021; de Chaisemartin and D’Haultfœuille, 2020; Strezhnev, 2018; Callaway and Sant’Anna, 2021; Borusyak, Jaravel and Spiess, 2024; Athey and Imbens, 2022). Most of this literature assumes staggered adoption, but the insights from that setting are still relevant when there are treatment reversals. In Figure 1, we present two simplified examples with staggered treatment adoption. Figure 1(a) represents outcome trajectories in line with standard TWFE assumptions, which not only include PT but also require that the treatment effect be immediate and unvarying across units and over time. In contrast, the right panel portrays a scenario where PT holds, but the constant treatment effect assumption is not met. Various decompositions by the aforementioned researchers reveal that even under PT, when treatments begin at different times (such as in staggered adoption) and treatment effects evolve over time and vary across units, the TWFE estimand is in general not a convex combination of the ITE for observations subjected to the treatment. The basic intuition behind this theoretical result is that TWFE models use post-treatment data from units who adopt treatment earlier in the panel as controls for those who adopt the treatment later (e.g., Goodman-Bacon, 2021). HTE-robust estimators capitalize on this insight by avoiding these “invalid” comparisons between two treated units in post-treatment periods. In the next section, we present a typology of HTE-robust estimators, along with an introduction to the estimators themselves.
A third limitation of the canonical TWFE model is its presumption of no temporal and spatial interference. In most uses of TWFE models, researchers assume that there are no spatial spillovers and that treatment effects occur contemporaneously, hence no anticipation or carryover effects. No anticipation effects means that future treatments do not affect today’s potential outcomes, while no carryover effects means that today’s treatment does not affect future potential outcomes: For any $i,t$,
(No anticipation effects) $Y_{i,t}(\mathbf{d}_{i})=Y_{i,t}(d_{i,1},d_{i,2},\ldots,d_{i,t})$
(No carryover effects) $Y_{i,t}(\mathbf{d}_{i})=Y_{i,t}(d_{i,t},d_{i,t+1}\ldots,d_{i,T})$
in which $\mathbf{d}_{i}=(d_{i,1},d_{i,2},\ldots,d_{i,T})$ is unit $i$’s vector of potential treatment conditions in all periods (from period $1$ to $T$) and $Y_{i,t}(\mathbf{d}_{i})$ is the potential outcome in period $t$ given $\mathbf{d}_{i}$. The TWFE model specified in Model (1) satisfies these two assumptions because $Y_{i,t}(\mathbf{d}_{i})=Y_{i,t}(d_{i,t})$. These assumptions are obviously very strong, but they are rarely questioned or tested in practice (Imai and Kim, 2019; Athey and Imbens, 2022; Wang, 2021). Although some recent methods permit arbitrary carryover effects in staggered adoption settings (Strezhnev, 2018; Callaway and Sant’Anna, 2021), these effects are not distinguishable from contemporaneous effects. [^4] See Section A.4. of the SM in Liu, Wang and Xu (2024) for more details. This limitation becomes more complex when treatment reversal is possible, as demonstrated in Figure 1. In Figure 1(b), data in the left panel are consistent with TWFE assumptions, while the right panel shows deviations from PT, constant treatment effect, and the absence of anticipation or carryover effects. What is concerning is that many real-world data resemble the problematic right figure rather than the ideal left one. Scholars have proposed methods to handle arbitrary carryover effects over a limited number of periods in the general setting (IKW, 2023; LWX, 2024). The challenge of addressing spatial spillover effects without strong structural assumptions still persists (Aronow, Samii and Wang, 2020; Wang, 2021), but its resolution is beyond the scope of this paper.
Causal Estimands.
[…]
The finest estimand is the ITE, $\tau_{i,t}=Y_{i,t}(1)-Y_{i,t}(0)$, of which there exists one for each observation $(i,t)$. [^6] This is without loss of generality when feedback and interference are excluded. In staggered DID designs, carryover effects are permissible. When potential outcomes are defined in terms of treatment history, $\tau_{i,t}$ is defined as $Y_{i,t}(1)-Y_{i,t}(\infty)$ where $Y_{i,t}(\infty)$ signifies the untreated potential outcome when unit $i$ never undergoes treatment. Most political science research, however, typically focuses on estimating a single summary statistic. Commonly, this is the ATT, which represents ITE averaged over all observations exposed to the treatment condition. In between these extremes of granularity and coarseness are time-varying dynamic treatment effects (DTE), which are across-unit averages of ITE at each time period relative to treatment adoption (e.g., all observations immediately proceeding treatment adoption). In the staggered adoption setting, we can further subdivide by cohort. We use $\delta_{l}$ ($\delta_{g,l}$) to denote the DTE $l$ periods after treatment adoption (for treatment cohort $g$), with $l=1$ to represent the period immediately after treatment adoption. [^7] Some of the authors we reference denote this first post-treatment period with $l=0$. $\delta_{g,l}$ is also what some authors refer to as a cohort average treatment effect on the treated (CATT) (Strezhnev, 2018; Sun and Abraham, 2021) or group-time average treatment effect (Callaway and Sant’Anna, 2021).
Each of the estimators we discuss can be used to estimate $\delta_{l}$. The outcome model analogous to TWFE for estimating DTE is a lags-and-leads specification. For simplicity, we first describe the staggered setting. Let $K_{i,t}=(t-E_{i,t}+1)$ be the number of periods until (when $K_{i,t}\leq 0$) or since unit $i$’s event time at time $t$ (e.g., $K_{i,t}=1$ if unit $i$ switches into treatment at time $t$). Consider a regression based on the following specification:
$Y_{i,t}=\alpha_{i}+\xi_{t}+X^{\prime}_{i,t}\beta+\sum_{\begin{subarray}{c}l=-a\\ l\neq 0\end{subarray}}^{b}\delta_{l}^{TWFE}\cdot\mathbf{1}\left\{K_{i,t}=l\right\}+\delta_{b+}^{TWFE}\mathbf{1}\left\{K_{i,t}>b\right\}\cdot D_{i,t}+\varepsilon_{i,t},\vspace{-1em}$ (2)
where $a$ and $b$ are the number of lag and lead terms (BJS, 2024). In the social science literature, the typical practice is to exclude $l=0$, which corresponds to the time period immediately before the transition into the treatment phase, and use it as a reference period (Roth, 2022). Conventionally, $\hat{\delta}^{TWFE}_{l}$ is interpreted as an estimate of $\delta_{l}$ or as a meaningful weighted average of pertinent ITE. Meanwhile, $\hat{\delta}^{TWFE}_{b+}$ is viewed as an estimate for the long-term effect.
3.  HTE-Robust Estimators
In this section, we offer a brief overview and comparison of several recently introduced HTE-robust estimators. We use the term “HTE” to refer to ITE that are nonidentical, i.e. $\tau_{i,t}\neq\tau_{j,s}$ for some $i,j,s,t$. HTE-robust estimators are defined as those that produce causally interpretable estimates under their respective identifying assumptions. For a more comprehensive discussion on all estimators mentioned in this section, please refer to the SM.
Summary of HTE-robust estimators.
Table 1 summarizes the estimators we discuss in this paper. The primary difference resides in the mechanics of their estimation strategies: There are methods based on canonical DIDs and methods based on imputation. We refer to the former as DID extensions and the latter as imputation methods. [^8] Liu, Sha and Zhang (2022) use a similar dichotomy to describe these estimators. DID extensions use DTE, estimated from local, $2\times 2$ DIDs between treated and control observations, as building blocks. Imputation methods use ITE, estimated as the difference between an imputed outcome under control and the observed outcome (under treatment), as building blocks. Imputation methods connect to TWFE through the outcome model, which is fit globally on all available controls, that they use to impute counterfactual outcomes. Different strategies also entail different assumptions. Each DID extension, for example, relies on a particular type of PT assumption, whereas imputation methods presuppose a TWFE model for untreated potential outcomes and require a zero mean for the error terms.
Another noteworthy difference lies in the settings in which these estimators are applicable: Some function exclusively in settings with staggered treatment adoption, while others can accommodate scenarios with treatment reversals. [^9] It is possible to use some methods designed for the staggered setting even when there are reversals by using only a subset of the data—in particular, by removing all observations that come after a treatment reversal—but this would identify a different effect than the ATT. In the latter setting, all estimators we discuss also require no anticipation and no or limited carryover effects. Furthermore, the estimators diverge in terms of (1) how they select untreated observations as controls for treated units, (2) how they incorporate pre-treatment or exogenous covariates, and (3) the choice of the reference period. We discuss these details further below and in Sections A.1. and A.4. of the SM.
Table 1: Summary of HTE-Robust Estimators
Type DID Extensions: uses $2\times 2$ DIDs as building blocks Imputation Methods
Setting Staggered: treatment reversals not allowed General: treatment reversals allowed
Research article Sun and Abraham (2021) Callaway and Sant’Anna (2021) Baker, Larcker and Wang (2022) de Chaisemartin and D’Haultfœuille (2020) IKW (2023) BJS (2024) LWX (2024)
Method known as IW CSDID StackedDID did_mulitple PanelMatch DIDimpute FEct
Key ID assumption Parallel trends Parallel trends Parallel trends Parallel trends Parallel trends Strict exogeneity Strict exogeneity
Finest estimand $\delta_{g,l}$ $\delta_{g,l}$ $\delta_{l}^{vw}$ $\delta_{l}$ $\delta_{l}$ $\tau_{i,t}$ $\tau_{i,t}$
Common aggregated estimand ATT ATT ATT ATT for switchers for $l$ periods ATT ATT
Comparison group Never-treated or last-treated Never-treated or not-yet-treated Never-treated Matched stable group (not-yet-treated) Matched stable group (not-yet-treated) Imputed Counterfactual (not-yet-treated) Imputed Counterfactual (not-yet-treated)
Reference period(s) Period 0 An arbitrary pre-treatment period Period 0 Untreated period Period 0 All pre-treatment periods All pre-treatment periods
Covariate adjustment Possible extension Outcome & propensity score modeling Outcome modeling Possible extension Refined matched set & outcome modeling Outcome modeling Outcome modeling
DID extensions.
DID extensions are all built from local, $2\times 2$ DID estimates—hence our choice of terminology. The overarching strategy for these estimators is to estimate the DTE, $\delta_{l}$ (or $\delta_{g,l}$ for each cohort $g$ in the staggered setting), for each period since the most recent initiation of treatment, $l$, using one or more valid $2\times 2$ DIDs. By “valid,” we mean that the DID includes (1) a pre-period and a post-period and (2) a treated group and a comparison group. The pre-period is such that all observations in both groups are in control, while the post-period is such that observations from the treated group are in treatment and those from the comparison group are in control. The choice of the comparison group is the primary distinction between estimators in this category. To obtain higher-level averages such as the ATT, we then average over our estimates of $\delta_{l}$ (or $\delta_{g,l}$), typically employing appropriate, convex weights.
We cover three estimators in this category that are appropriate only for the staggered setting. Sun and Abraham (2021) propose an interaction-weighted (IW) estimator, which is a weighted average of $\delta_{g,l}$ estimates obtained from a TWFE regression with cohort dummies fully interacted with indicators of relative time to the onset of treatment. They demonstrate that each resulting estimate of $\delta_{g,l}$ can be characterized as a difference in the change in average outcome from a fixed pre-period $s<g$ to a post-period $l$ periods since $g$ between the treated cohort $g$ and the comparison cohort(s) in some set $\mathcal{C}$. The authors recommend using $\mathcal{C}={\sup_{i}E_{i}}$, which is either the never-treated cohort or, if no such cohort exists, the last-treated cohort. By default, IW uses $l=0$ as the reference period and can accommodate covariates in the TWFE regression.
Employing the same general approach, Callaway and Sant’Anna (2021) propose two estimators, one of which uses never-treated units ($\hat{\delta}^{CS-dr}_{nev}$) and the other not-yet-treated units ($\hat{\delta}^{CS-dr}_{ny}$) as the comparison group. We label these estimators collectively as CSDID. Note that $\hat{\delta}^{CS-dr}_{nev}$ uses the same comparison group as IW when a never-treated cohort exists, [^10] This equivalence holds when there are no missing data. Otherwise, IW from the saturated regression differs from one that directly estimates local DIDs, including the never-treated version of CSDID. These estimates are typically close but can differ substantially, as in Kuipers and Sahn (2023). whereas $\hat{\delta}^{CS-dr}_{ny}$ uses all untreated observations of not only never-treated units but also later adopters as controls for earlier adopters. Besides the choice of comparison cohort, CSDID estimators differ from IW in that they allow users to condition on pre-treatment covariates using both an explicit outcome model and inverse probability weighting (IPW) simultaneously, with at least one needing to be correct for the estimator to be consistent. By default, both IW and CSDID use the period immediately before the treatment’s onset as the reference period for estimating the ATT.
Stacked DID is another related estimator sometimes used to address HTE concerns. As described by Baker, Larcker and Wang (2022), it involves creating separate sub-datasets for each treated cohort by combining data from that cohort (around treatment adoption) and data from the never-treated cohort from the same periods. These cohort-specific datasets are then “stacked” to form a single dataset. An event study regression akin to Model (2) with the addition of sub-dataset specific unit and time dummies is then run. This method uses the same comparison group as IW and the never-treated version of CSDID without covariates, but stacked DID estimates a single DTE for a given relative period rather than separate estimates for each cohort. Essentially, stacked DID is a special case of IW that uses immutable weights selected by OLS. We denote the corresponding estimand $\delta_{l}^{vw}$ to reflect the fact that it is variance-weighted. These weights will be convex but in general are not proportional to cohort sizes. Thus, while stacked DID avoids the “negative weighting” problem and meets our criteria for HTE-robustness, its estimands are not the same as $\delta_{l}$ or the ATT. This, however, is arguably a minor issue related to external validity, as the estimands can be interpreted as average treatment effects for populations with a different distribution over ITE than the treated observations.
In settings with treatment reversals, separate groups of researchers have converged on the same strategy for choosing a comparison group: matching treated and control observations that belong to units with identical treatment histories. IKW (2023) suggest one such estimator, PanelMatch, which begins by constructing a “matched set” for each observation $(i,t)$ such that unit $i$ transitions into treatment at time $t$. This matched set includes units that both (1) are not under treatment at time $t$ and (2) share the same treatment history as $i$ for a fixed number of periods leading up to the treatment onset. For each treated observation $(i,t)$ and for every post-period $t+l-1$ such that unit $i$ is still under treatment, it then estimates a local DID using the same pre-period $s<t$. The treatment “group” comprises solely of $(i,t)$, and the members of the matched set for $(i,t)$ that are still under control during $t+l-1$ serve as the comparison group. To obtain $\delta_{l}$ for a given $l$, it then averages over the corresponding local DID estimates from all treated observations. IKW (2023) propose incorporating covariates by “refining” matched sets and use $l=0$ as the reference period.
[…]
All DID extensions are built using local, $2\times 2$ DIDs, and their assumptions reflect this. Specifically, they each rely on a form of the PT assumption—that is, the expected changes in untreated potential outcomes from one period to the next are equal between the treated and the chosen comparison groups. We defer readers to the SM for a fuller account of each method’s assumptions.
The imputation method.
Imputation estimators do not explicitly estimate local DIDs. Instead, they take the difference between the observed outcome and an imputed counterfactual outcome for each treated observation. The connection to the TWFE model is in the functional form assumption used to impute counterfactual outcomes. Specifically, an imputation estimator first fits a parametric model for the potential outcome under control $Y_{i,t}(0)$—in our case, $Y_{i,t}(0)=X_{i,t}{{}^{\prime}}\beta+\alpha_{i}+\xi_{t}+\varepsilon_{i,t}$—using only control observations $\{(i,t):D_{i,t}=0\}$. It is also through this outcome model that one can adjust for time-varying covariates. Then, it imputes $Y_{i,t}(0)$ for all treated observations $\{(i,t):D_{i,t}=1\}$ using the estimated parameters. Finally, it estimates the ITE, $\tau_{i,t}$, for each treated observation $(i,t)$ by calculating the difference between the observation’s observed outcome $Y_{i,t}=Y_{i,t}(1)$ and its imputed counterfactual outcome $\hat{Y}_{i,t}(0)$. Inference for the estimated $\hat{\tau}_{i,t}$ is possible, although uncertainty estimates need to be adjusted to account for the presence of idiosyncratic errors (e.g., Bai and Ng, 2021). BJS (2024) and LWX (2024) each propose estimators in this category. Each paper proposes a more general framework that nests many models, including TWFE. The latter also introduces several specific imputation estimators, one of which uses the TWFE model, and the authors refer to the resulting estimator as the fixed effect counterfactual estimator, or FEct.
Although DID extensions and imputation methods rely on slightly different identification assumptions, these assumptions usually lead to similar observable implications. Researchers commonly use the presence or absence of pretrends to judge how plausible the PT assumption is. In the classic two-group setting, if there are data from multiple pre-treatment periods, researchers can plot the time series of average outcomes of each group and visually inspect whether they indeed trend together. The intuition is that if PT holds and the outcome trends of the treated and control groups are indeed parallel in pre-treatment periods when $Y(0)$’s are observed for all units, then it is plausible that PT also holds in the post-treatment periods, when $Y(0)$’s are no longer observable for units in the treatment group. Conversely, differential trends in the pre-treatment periods should make us suspicious of PT. In more complex settings or where we wish to control for observed confounders, researchers often use estimates of the dynamic effects before and after the onset of treatment, $\delta_{l}$, to construct so-called “event-study plots” to judge the presence of pretrends. If PT holds, then pre-treatment DTE estimates should be around zero. We provide a more thorough discussion and an example of the event-study plot in the next section when we introduce our procedure.
Choice of estimators.
In general, we believe the credibility of identifying assumptions is more important than the choice of estimator. After all, in the staggered setting, when assumptions hold, IW, CSDID, PanelMatch, DIDM, and the imputation estimator all converge to the same or a similar estimand. However, there are a few reasons to favor imputation estimators. First, imputation estimators can handle complex settings, including those with treatment reversal—which account for over half of the studies in our sample—and can accommodate time-varying covariates, additional fixed effects, and unit- or group-specific time trends commonly seen in social science research. Second, imputation estimators are the most efficient under homoskedastic errors (BJS, 2024). [^12] In Section A.4.2 of the SM, we demonstrate that the imputation estimator is generally more efficient, based on data from our sample. Moreover, by using the average of all pre-treatment periods as the reference point rather than a single pre-period, as in DID extensions, they provide greater power in hypothesis testing for pretrends (see Section A.1.6 of the SM for Monte Carlo evidence). The main drawback of imputation estimators is that their current implementations (either FEct or DID_impute) does not allow for automated adjustment of time-invariant covariates, an advantage offered by CSDID and PanelMatch. Adjusting for pre-treatment characteristics can improve credibility of research, as conditional PT may be more plausible than the unconditional one (Sant’Anna and Zhao, 2020).
4.  Data and Procedure
Next, we assess the robustness of empirical findings from causal panel analyses in political science and compare results obtained using the different methods we have discussed. We will explain our sample selection rules, describe standard practices in the field, and outline our reanalysis approach. Readers can find a more detailed explanation of our sample selection criteria and replication and reanalysis procedure in Sections A.3. and A.4. of the SM.
Data.
Our replication sample comprises articles from three leading political science journals, the APSR, AJPS, and JOP, published over a recent seven-year span from 2017 to 2023. We initially include all studies, including both long and short articles, that employ panel data analyses with a binary treatment as a crucial component of their causal argument, resulting in a total of 102 articles. After a careful review of each of these articles, we find that 64 articles use a TWFE model similar to Model (1). We then attempt to replicate the main results of these 64 articles and are successful in 49 cases (76.6%). Though a significant proportion of studies failed to replicate, we note that the success rate is still higher than that of Hainmueller, Mummolo and Xu (2019) at 55% and is comparable to that of Lal et al. (2024), which stands at 67%. We credit this to the new replicability standards set by journals. A detailed explanation of how we select the “main model” is provided in Section A.3. of the SM. Table 2 depicts the distribution of successful replications, along with reasons for replication failures, across the various journals.
Table 2: Sample Selection and Replicability of Qualified Articles
TWFE Incomplete Replication Success
Journal All (attempted) data error Replicable Rate%
APSR 22 13 2 1 10 76.9
AJPS 31 21 3 3 15 71.4
JOP 49 30 6 0 24 80.0
Total 102 64 11 4 49 76.6
Settings and common practices.
Table 3 presents an overview of the standard practices and settings in the articles that we successfully replicated. The vast majority of studies in our sample (93.9%) use the DID design/method/approach to justify the use of the TWFE model, while the remaining studies advocate for the model’s ability to exploit “within” variations in the data. Out of the 49 articles, nine (18.4%) employ a classic DID design, which includes two-group, two-period designs (3 articles) and multi-period DID designs (6 articles). 13 articles (26.5%) use a staggered (but not classic) DID design, while the remaining 27 articles (55.1%) fall into the “general” category, meaning they allow for treatment reversals. Except for four articles, all studies have a continuous outcome of interest. Most studies adopt cluster-robust SEs or panel-corrected SEs (Beck and Katz, 1995), while eight articles apply bootstrap procedures for estimating uncertainties. A subset of authors explore alternative model specifications by adding lagged dependent variables (8 articles), unit-specific linear time trends (15 articles), and higher-than-unit-level time trends (5 articles). Notably, 30 studies use some type of visual inspection—either average outcomes over time, event-study plots, or both—to evaluate the plausibility of PT. Four studies published in 2023 (33%) employ HTE-robust estimators, compared to none before 2023, indicating rapid adoption of these methods. Of these, two use CSDID, one PanelMatch, and one the imputation estimator. Roughly 40% of the studies in our sample do not include any visual inspection, such as an event-study plot, in the main text or appendices.
Table 3: Settings and Common Practice
Motivations for TWFE Variance Estimator
“Difference-in-differences” 46 93.9% Cluster-robust SE or PCSE 48 98.0%
“Within” variations 3 6.1% Cluster-bootstrapped procedures 8 16.3%
Treatment Setting Variants in specifications
Classic 2$\times$2 DID 3 6.1% Lagged dependent variables 8 16.3%
Classic multi-period DID 6 12.2% Higher-than-Unit-level time trends 5 10.2%
Staggered DID 13 26.5% Unit-specific linear time trends 15 30.6%
General 27 55.1%
Data visualization
Outcome Variable Group average outcomes 19 38.8%
Continuous 44 89.8% Event-study plots 23 46.9%
Binary 5 10.2% Neither 19 38.8%
Procedure.
We use data from Grumbach and Sahn (2020) to illustrate our process for replication and reanalysis. The study investigates the influence on coethnic mobilization by minority candidates during US congressional elections. To simplify our analysis, we focus on the impact of the presence of an Asian candidate on the proportion of general election contributions from Asian donors. To begin, we aim to understand the research setting and data structure. We visualize the patterns of treatment and outcome variables using plots, which are shown in the SM. In this application, treatment reversals clearly take place. Some data is missing (due to redistricting), but the issue does not seem to be severe. We record important details such as the number of observations, units, and time periods, the type of variance estimator, and other specifics of the main model (not shown here). Next, we replicate the main finding, employing both the original variance estimator and a cluster-bootstrap procedure.
We then re-estimate the ATT and DTE using estimators discussed in Section 3. For staggered adoption treatment cases, we apply seven estimators: TWFE (with always treated units removed for easier comparisons with other estimators), the imputation estimator (FEct), PanelMatch, DIDM, StackedDID, IW, and CSDID (both not-yet-treated and never-treated versions). For applications with treatment reversals like Grumbach and Sahn (2020), we use the first two estimators only. [^13] When there are treatment reversals, the current release of the DIDM software implements a different estimator than the one described in de Chaisemartin and D’Haultfœuille (2020). In particular, it only estimates the contemporaneous effect at the first onset of treatment as if there are no treatment reversals. Therefore, we only apply DIDM to the staggered cases. The comparison between the TWFE estimate and the other estimates sheds light on whether original findings are sensitive to relaxing the constant treatment effect assumption. Figures 2(a)-(d) show the results from this example. The similarity between estimates for the ATT in (a) suggests that the original finding is robust to the choice of estimators. The event-study plots from HTE-robust estimators in (c) and (d) are broadly consistent with the event-study plot from TWFE in (b).
Figure 2: Reanalysis of Grumbach and Sahn (2020)
(a) Treatment effect estimates
(b) Dynamic: TWFE
(c) Dynamic: PannelMatch
(d) Dynamic: imputation
(e) Placebo test
(f) Test for carryover effects
Note: Reanalysis of data from Grumbach and Sahn (2020). Subfigures (a): Treatment effect or ATT estimates from multiple methods; subfigures (b)-(d): event-study plots using TWFE, PanelMatch, and the imputation estimator (FEct); Subfigure (e)-(f): results from the placebo test and test for carryover effects using the imputation estimator—the blue points in (g) and red points in (h) represent the holdout periods in the respective tests. CIs in all subfigures—excepted for the reported estimate in (a)—are produced by bootstrap percentile methods. The robust confident set for the ATT ($\bar{M}=0.5$) is shown in pink in (e).
Next, we conduct diagnostic tests based on the imputation estimator, including the $F$ test and the placebo test, to further assess the plausibility of PT and, in applications with treatment reversal, the no-carryover-effect assumption. We use the imputation estimator because it is applicable across all studies in our replication sample, can incorporate time-varying covariates, and remains highly efficient. Figures 2(d), (e) and (f) show the results from the $F$ test, placebo test, and test for no carryover effects on our running example, respectively. Both a visual inspection and the formal tests suggest that PT and no-carryover-effect assumptions are quite plausible.
Finally, we compute robust confidence sets (CSs) proposed by Rambachan and Roth (2023), a partial identification approach that avoids the conditional inference problem by accounting for potential PT violations when testing the null hypothesis of no treatment effect in the post-treatment period. Specifically, we use relative magnitude (RM) restrictions, which allow for post-treatment PT violations that are at most $\bar{M}$ times the size of the maximum violation estimated in pre-treatment placebo periods. In other words, the differential trend between any two consecutive post-treatment periods are assumed to be no greater than $\bar{M}$ times the maximum difference between any two placebo periods. Using the placebo periods ensures that the benchmark pre-treatment estimates are obtained in the same way as post-treatment ATT estimates, avoiding potential asymmetry in testing and treatment effect estimation (Roth, 2024). If we first assume we can decompose each estimator for the DTE $\beta_{t}$ into the actual DTE $\tau_{t}$ and a trend component $\delta_{t}$, $\beta_{t}=\tau_{t}+\delta_{t}$, then the RM assumption requires that for all periods relative to treatment $t$,
$|\delta_{t+1}-\delta_{t}|\leq\bar{M}\cdot\max_{s\in\mathcal{P}\backslash\{0\}}\{\delta_{s+1}-\delta_{s}\}$.
in which $\mathcal{P}$ is the set of placebo periods excluding the period immediately preceding treatment adoption. In our case, we use $\mathcal{P}=\{-2,-1\}$, meaning the maximum trend is given by $\max\{\delta_{0}-\delta_{-1},\delta_{-1}-\delta_{-2}\}$. Robust CSs are uniformly valid for the partially identified treatment effects under the RM restriction. Rambachan and Roth (2023) suggest $\bar{M}=1$ as a “natural benchmark” when the number of pre-treatment periods used for testing is comparable to the number of post-treatment periods, assuming that any PT violations are no worse than the observed pretrend (p. 2653). [^14] In our case, the number of post-treatment periods typically exceeds the number of placebo periods, resulting in a laxer standard, as the maximum is non-decreasing with the number of pre-treatment periods. For each study, we compute robust CSs with $\bar{M}=0.5$ to apply a lenient benchmark. For each study, we compute the robust CSs with $\bar{M}=0.5$, erring on the lenient side. We also conduct a sensitivity analysis for each such study, obtaining the “breakdown value” of $\bar{M}$, denoted as $\tilde{M}$, at which the robust CS first includes zero. Figure 2(e) marks the robust CS for the ATT estimated using the imputation estimator in pink. In this example, the robust CS excludes zero, and the breakdown value from the sensitivity analysis is $\tilde{M}=3.0$, suggesting that the estimated coethnic mobilization effect is robust to potential PT violations under the RM restriction.
Overall, the results from Grumbach and Sahn (2020) appear highly robust, regardless of the choice of point and variance estimators. The PT and no-carryover-effect assumption seem plausible. The study also has sufficient power to distinguish the ATT from zero, even under potential PT violations.
5.  Systematic Assessment
We perform the replication and reanalysis procedure described above for all 49 articles in our sample. This section offers a summary of our findings, with complete results for each article available in the SM. We organize our results around two main questions: (1) Are existing empirical findings based on TWFE models robust to HTE-robust estimators? (2) Is the PT assumption plausible, and do original findings remain robust to mild PT violations informed by pretrends? We also discuss other issues observed in the replicated articles, including the presence of carryover effects and sensitivity to model specifications.
HTE-robust estimators yield qualitatively similar but highly variable findings.
To examine the impact of the weighting problem caused by HTE associated with TWFE models, we first compare the estimates obtained from the imputation estimator, FEct, for all studies to those originally reported. The dissimilarity between the original and FEct estimates proxies the extent to which the negative weighting problem is consequential in practice. We choose the imputation estimator for the reason mentioned in Section 3. Most importantly, it is applicable to all studies in our sample, including those with treatment reversals. Figure 3 plots the comparison. The horizontal axis represents the originally reported TWFE estimates, and the vertical axis represents estimates from the imputation estimator, both normalized using the same originally reported SE. If the point estimates are identical, then the corresponding point should lie exactly on the 45-degree line. Red triangles represent studies where the imputation estimates are statistically insignificant at the 5% level, based on cluster-bootstrapped SEs.
Figure 3: TWFE vs. The Imputation Estimator: All Cases
Note: The above figure compares reported TWFE coefficients with imputation method (FEct) estimates. Both estimates for each application are normalized by the same reported TWFE SE. Fouirnaies and Hall (2018) and Hall and Yoder (2022) are close to the 45-degree line but are not included in the figure as their TWFE $z$-scores exceed 15. Black circles (red triangles) represent studies whose imputation method estimates for the ATT are statistically significant (insignificant) at the 5% level, based on cluster-bootstrapped SEs. The top-left (bottom-right) corners display histograms of the ratio of point (SE) estimates based on the imputation method and TWFE. These plots show that changes in point estimates, combined with the efficiency loss from using the imputation method, contribute to the loss of statistical significance in some studies.
We observe several patterns. First, TWFE coefficients are statistically significant at the 5% level in all but one study, and the absolute values of $z$-scores for a significant minority of studies cluster around $1.96$, indicating possibly a file-drawer problem and potential publication bias. Second, the points largely follow the 45-degree line, with the imputation estimates always having the same sign as the original estimates. This suggests that while scenarios where accounting for HTE completely reverses the empirical findings are theoretically possible, they are rare.
However, results sometimes deviate significantly. In the top left corner of Figure 3, we plot the histogram of the ratio of imputation to TWFE estimates. Although the mean and median of the ratio are close to one, at the extremes, we observe imputation estimates as small as one-fourth or as large as more than double the TWFE estimates. The most consequential deviations occur in studies that were originally near the margins of statistical significance. Additionally, we plot the ratio of SE estimates from the imputation method to TWFE in the bottom right corner of Figure 3. The median is 1.1, meaning that in the majority of cases the SE estimate from the imputation method is at least 10% larger than the SE from TWFE. The mean is 1.23, and the distribution is right-skewed; in the extreme, the ratio was almost as high as three. Combined, drops in point estimates and increases in uncertainty lead to the third pattern: When switching from TWFE to the imputation estimator, the number of studies that are statistically insignificant at the 5% level increases from one to 12 (24%).
If we restrict our attention to eleven studies with staggered treatments, we can broaden our comparison set to include more HTE-robust estimators discussed earlier. [^15] Kogan (2021) and Magaloni, Franco-Vivanco and Melo (2020) are excluded because the original specifications include additional time trends, which are not supported by HTE-robust estimators except the imputation estimator. Figure 4 visually compares the points estimates (left panel) and $z$-scores (right panel). In the left panel, all point estimates are normalized by the same reported TWFE SE for each study. In all but three studies, the estimates from all HTE-robust estimators share the same sign, though there is a noticeable amount of variation in the estimated effect size. [^16] Although a sign change is observed in Clayton and Zetterberg (2018) when using CSDID, IW, and DIDM, these estimates are negligibly small and statistically insignificant. Similarly, the not-yet-treated version of CSDID is the opposite sign but miniscule and statistically insignificant in Bischof and Wagner (2019). The estimates from IW and CSDID are also of the opposite sign for Kuipers and Sahn (2023), and they are of a much larger magnitude. The never-treated version of CSDID is also statistically significant.
Figure 4: Comparison of Estimates: The Staggered Setting
Refer to captionRefer to caption
Note: The above figures compare reported TWFE coefficients with estimates from various alternative estimators. In the left panel, all eight estimates for each application are normalized by the same reported SE to highlight changes resulting from the use of alternative estimators. In the right panel, the estimates are divided by their respective bootstrapped SEs. To facilitate visualization, we multiply all estimates by the sign of the reported coefficient. In both figures, black (red) circles represent estimates that would be statistically significant (insignificant) at the 5% level, assuming they were treated as $z$-scores. The normalized CSDID (never treated) estimate for Kuipers and Sahn (2023), -5.36, falls out of plotting area and is therefore not shown in the left panel. Kogan (2021) and Magaloni, Franco-Vivanco and Melo (2020) are excluded because the authors’ original TWFE specifications include unit-specific linear time trends, which are not supported by the HTE-robust estimators except the imputation estimator. Some estimates are missing because of too few never-treated units. PanelMatch is excluded because it targets a different estimand.
As in Figure 3, TWFE does not appear to be systematically upward or downward biased compared to HTE-robust estimators. Another observation that carries over is that HTE-robust estimators generally require more power to reject the null hypothesis of no effect. In five of the eleven studies, at least four HTE-robust estimates per study are statistically insignificant. The left panel shows that the changes in point estimate alone are often sufficient to render the results statistically insignificant. The comparison of $z$-scores in the right panel highlights that increased uncertainties can be substantial. Combined with earlier evidence, these findings from the staggered cases suggest that HTE bias in TWFE is empirically significant and warrants careful consideration by researchers.
It is worth noting that while only a small fraction of studies in our sample (five articles, 13.5%) employ a bootstrap procedure to estimate SEs or CIs, the widely practiced cluster-robust SE typically performs adequately. This is because the number of units (clusters) is generally large, with a median of 317. However, 10 studies have fewer than 50 units; among them, two studies that were significant at the 5% level using cluster-robust SEs fell below this threshold when using cluster-bootstrapped SEs, both of which were already marginally significant. We provide comparisons of reported, cluster-robust, and cluster-bootstrapped SEs in the SM.
Figure 5: Estimated Dynamic Treatment Effects
Beazer and Reuter (2022)
ATT: 24.38 (8.80) Bischof and Wagner (2019)
ATT: 0.08 (0.07) Bisgaard and Slothuus (2018)
ATT: -0.05 (0.02) Blair, Christensen and Wirtschafter (2022)
ATT: -0.50 (0.39) Bokobza et al. (2022)
ATT: 0.11 (0.03)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Caughey, Warshaw and Xu (2017)
ATT: 0.01 (0.005) Christensen and Garfias (2021)
ATT: 0.04 (0.04) Clarke (2020)
ATT: 0.08 (0.03) Clayton and Zetterberg (2018)
ATT: 0.19 (0.44) Cox and Dincecco (2021)
ATT: 0.44 (0.11)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Dahlström and Holmgren (2023)
ATT: -1.67 (0.62) Dipoppa, Grossman and Zonszein (2023)
ATT: 0.001 (0.003) Distelhorst and Locke (2018)
ATT: 0.15 (0.05) Eckhouse (2022)
ATT: 0.01 (0.01) Esberg and Siegel (2023)
ATT: 0.01 (0.01)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Fouirnaies (2018)
ATT: 0.87 (0.11) Fouirnaies and Hall (2018)
ATT: 0.29 (0.02) Fouirnaies and Hall (2022)
ATT: -0.26 (0.03) Fresh (2018)
ATT: 0.17 (0.06) Garfias (2019)
ATT: 0.07 (0.03)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Grumbach (2023)
ATT: -0.57 (0.15) Grumbach and Sahn (2020)
ATT: 0.13 (0.03) Grumbach and Hill (2022)
ATT: 0.033 (0.018) Hainmueller and Hangartner (2019)
ATT: 1.51 (0.21) Hankinson and Magazinnik (2023)
ATT: -0.51 (0.44)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Estimated Dynamic Treatment Effects (Cont.)
Hall and Yoder (2022)
ATT: 0.06 (0.003) Hirano et al. (2022)
ATT: 0.56 (0.18) Jiang (2018)
ATT: -0.87 (0.20) Kilborn and Vishwanath (2022)
ATT: 0.12 (0.04) Kogan (2021)
ATT: 0.45 (0.39)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Kroeger and Silfa (2023)
ATT: 0.03 (0.01) Kuipers and Sahn (2023)
ATT: 0.05 (0.02) Liao (2023)
ATT: -0.04 (0.01) Magaloni, Franco-Vivanco and Melo (2020)
ATT: -2.28 (0.96) Marsh (2023)
ATT: -0.01 (0.004)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Latura and Weeks (2023)
ATT: 0.03 (0.01) Paglayan (2022)
ATT: 6.45 (2.09) Payson (2020a)
ATT: 0.04 (0.02) Payson (2020b)
ATT: 13.04 (8.44) Pierskalla and Sacks (2018)
ATT: -0.07 (0.04)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ravanilla, Sexton and Haim (2022)
ATT: 0.36 (0.18); Sanford (2023)
ATT: -2.35 (0.17) Schafer et al. (2022)
ATT: -0.02 (0.003) Schubiger (2021)
ATT: 0.05 (0.01) Schuit and Rogowski (2017)
ATT: -0.16 (0.04)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Trounstine (2020)
ATT: -0.06 (0.01) Weschle (2021)
ATT: 1.11 (0.72) Zhang et al. (2021)
ATT: 0.11 (0.06)
Refer to caption Refer to caption Refer to caption Note: We report the estimated ATT and corresponding bootstrap SEs (in parentheses) using the imputation estimator, FEct. We provide the $p$-values from the $F$ test on the top-left corner in each figure; rejecting the null in this test indicates potential PT violations. Note that these tests are infeasible for four cases with only one pretreatment period.
Relaxing PT renders most studies unable to reject the null.
Although the recent methodological literature has heavily focused on HTE, PT violations—long recognized as a potential pitfall—remain a primary concern in practice. In Figure 5, we present the event-study plot based on estimates from the imputation estimator for each study in our sample. We also report the ATT estimates and their bootstrapped SEs, as well as the $p$-values of the $F$ tests used to assess pretrends. Due to space limitations, we present the event-study plots from other estimators, as well as results from the placebo tests, robust CSs, and sensitivity analyses, in the SM.
What is concerning is that, in at least six studies (12%), the PT assumption seems highly implausible. In these studies, the DTE estimates in the pre-treatment periods deviate substantially from zero compared with those in the post-treatment period, and the $F$ test rejects the null. [^17] There are cases where the $F$ test rejects, but we do not consider them highly problematic because, with a large sample size, a confounder inconsequential to the ATT estimate can still produce a small $p$-value in the $F$ test. This is why the sensitivity analysis approach is particularly useful. For the remaining studies, the CIs of pre-treatment estimates mostly cover zero, and the $F$ test and placebo test do not reject. However, this could be simply due to a lack of statistical power. Therefore, to assess the robustness of the findings, we need additional tools that simultaneously account for both the estimated pretrend and statistical power. The sensitivity analysis proposed by Rambachan and Roth (2023) does exactly that.
Figure 5: Allowing PT violations with Robust Confidence Sets
(a) Robust infernece under PT violations
(b) Distribution of breakdown threhsold $\tilde{M}$
Note: The above figures present findings from the sensitivity analysis. The sample consists of 42 studies with more than three pre-treatment periods, allowing for such an analysis. Subfigure (a) displays the $p$-values of partially identified ATT estimates using the imputation method under restricted RM PT violations with $\bar{M}=0.5$, compared to reported $p$-values for TWFE coefficients assuming PT ($\bar{M}=0$). A square root scale is used to facilitate visualization. Black (red) triangles represent studies that are statistically significant (insignificant) at the 5% level when using the imputation method under PT. Subfigure (b) shows a histogram of $\tilde{M}$, the breakdown values of $\bar{M}$; the pink bar represents 11 studies whose ATT estimates are statistically insignificant at the 5% level when using the imputation method.
We conduct a sensitivity analysis for 42 studies with at least three pre-treatment periods, for which such a test is possible. In the vast majority of these studies, when accounting for potential PT violations—even very mild ones based on estimated pretrends in the placebo periods—we cannot conclude that the ATT is statistically significantly different from zero. In other words, the ATT estimates are not large enough to be distinguished from realistic PT violations, or from estimation and sampling errors. Figure 5(a) shows that the null of no effect is rejected at the 5% level in only 11 (26%) among the 42 studies with $\bar{M}=0.5$. Figure 5(b) displays the distribution of breakdown values $\tilde{M}$. Among the 42 studies, the median is 0.15 and the mean is 0.52. Focusing only on the 29 studies that remain statistically significant at the 5% level with the imputation estimator, the median and mean are still as low as 0.35 and 0.76, respectively. Although using HTE-robust estimators with robust CSs does not necessarily “overturn” results in most of the studies we reanalyzed, it is clear that the vast majority of studies lack the power to discern the ATT from zero under very mild PT violations of realistic magnitudes, as informed by pretrends.
Other issues.
Our reanalysis highlights several additional issues. First, the presence of missing values is widespread. Although most methodological work presumes balanced panels without missing data, in reality, many empirical studies encounter varying degrees of data missingness. Substantial differences in results for estimators that are numerically identical in balanced panels suggest that such violations may have substantial implications. We plot patterns of treatment status for each study in the SM. Based on these plots, we also see that in some studies the pattern of missingness is either seemingly nonrandom or extremely prevalent, which weakens our confidence in the respective empirical findings.
Second, we perform tests for carryover effects for all studies that feature treatment reversals. If this test is rejected, it suggests that the treatment effects persist beyond the treatment periods. Among 27 articles, 5 reject the null hypothesis of no carryover effects at 5%. Part of the concern is that imputation methods and DID extensions will use control observations from previously treated units to fit the potential outcome model or as comparisons for treated observations, and if there are carryover effects, then the comparisons will become tainted. LWX (2024) note that the presence of carryover effects for a limited number of periods is less concerning, as researchers can recode treatment to persist for some time after a unit transitions out of treatment. Despite its prevalence, we observe that carryover effects do not substantially alter the findings in most studies. Specifically, when removing two post-treatment periods for studies that reject the null of no carryover effects, the ATT estimates remain similar in magnitude. Additionally, results initially deemed statistically significant continue to maintain their statistical significance (Figure A11 in the SM). Nevertheless, we recommend that researchers make it a practice to check for potential carryover effects, considering the low cost of conducting such tests and adjustments.
Finally, many studies also exhibit sensitivity to model specifications. Some studies that we exclude from our sample employ one-way FE or FE at a level different from that at which treatment is assigned. Many of these findings do not hold when we reanalyze them using a TWFE model. We should clarify that this does not imply that the original results are wrong; rather, these models implicitly operate under different identifying assumptions, and there is substantial variation in how much consideration authors give to this point. Some studies do not provide a rationale for their choice to use one-way fixed effects, while others explicitly outline the type of unobserved confounders they intend to control for. In one instance, the authors inaccurately label their specification as a DID research design. The TWFE and DID estimators are generally not equivalent, and we emphasize that this difference becomes even more pronounced when fixed effects are not assigned at the level of treatment. In such cases, a TWFE specification does not correspond to even a broadly defined DID research design.
Summary.
In Table 4 below, we summarize the main findings of our reanalysis. The numbers represent the proportion of studies in a given journal or across all journals that fall under the respective category (hence, a lower number is better). The first two columns relate to the ratio of the imputation estimator to the TWFE estimator, which proxies the consequences of the weighting problem due to HTE. Across all journals, this ratio is less than 0.8 in 18% of studies and less than 0.5 in 6% of studies, indicating that using an HTE-robust estimator does sometimes have a substantively significant impact and is important to use at least as a robustness check.
Table 4: Summary of Findings
Consequence of the weigthing problem Potential PT violations Insufficient power w/. relaxed assumptions
(1) (2) (3) (4) (5) (6) (7)
Imputation-TWFE Imputation-TWFE $F$ test Placebo test TWFE not Imputation not Imputation not reject
Journal Ratio $<0.8$ Ratio $<0.5$ reject null reject null reject null reject null null w/. $\bar{M}=0.5$
$n$ 49 49 44 44 49 49 42
APSR (16) 0.00 0.00 0.40 0.40 0.00 0.10 0.70
AJPS (10) 0.11 0.00 0.21 0.29 0.00 0.12 0.69
JOP (23) 0.30 0.13 0.30 0.15 0.04 0.39 0.79
All (49) 0.18 0.06 0.30 0.25 0.02 0.24 0.74
Note: Entries are proportion of studies satisfying each set of conditions. A null is deemed being rejected if $p<0.05$. In columns (1)–(2), “ratio” refers to the ratio of the Imputation (FEct) estimate to TWFE estimate for the average effect. Five studies with a single pre-treatment period are not included in the summary statistics in columns (3)–(4). Seven studies are not included in column (7) because of too few pre-treatment periods.
Columns (3) and (4) indicate the proportion of studies with clear evidence of PT violations. Specifically, the null hypothesis is rejected at the 5% significance level for the $F$ test or the placebo test. Across all journals, around a quarter of studies reject the null hypothesis using the $F$ test, with slightly fewer rejecting when using the placebo test due to the loss in power from excluding data from the placebo periods. Note, though, that failure to reject can result from insufficient power and is not sufficient to suggest that the PT holds, but it is a modest standard that credible studies should be able to clear. [^18] On the other hand, it is also important to note that rejection does not always mean a result is invalid. We reiterate that pre-trends can reflect but are not the same as the trends in the statement of the PT. Moreover, a test can sometimes be too powerful and reject due to negligibly small pre-trends. It is thus also important to consider other evidence, such as the results from the robust CSs, that can vindicate such studies. Columns (5) and (6) display the proportion of studies in which the null hypothesis of no effect is not rejected using the TWFE and imputation estimators, respectively. When TWFE is employed, this happens in only one study (2%); however, when the imputation estimator is used, this number increases to 24%, suggesting many studies in our sample are potentially underpowered. Moreover, column (7) shows that in 74% of the studies, the robust CSs for the ATT include zero with $\bar{M}=0.5$. This further underscores the fragility of results in the existing literature when we allow even very mild PT violations.
Our reanalysis is not meant to criticize existing studies, many of which were conducted before recent methodological advances. In fact, we have observed rapid adoption of these new methods and greater statistical power in more recently published studies. However, we aim to emphasize two points: (1) many studies do not adequately assess the plausibility of PT, the key identifying assumption of their research designs; and (2) given recent methodological developments, causal panel analysis under the PT assumption requires significantly more power to account for HTE and potential PT violations than previously believed.
6.  Recommendations
Based on the findings of the reanalysis, we provide the following recommendations.
Research design is the key.
An important component of a strong research design is a clear understanding of how treatment is assigned. The PT assumption, which both TWFE models and most modern methods rely on, is silent on the assignment mechanism. As a result, many researchers assume that these methods can be applied when the assignment mechanism is unknown and that the absence of a pretrend is sufficient to make PT credible. Another perspective suggests that (quasi-)randomization is required for PT to be credible (e.g., Kahn-Lang and Lang, 2020). Our view aligns more closely with the latter, but is less stringent. We argue that for PT to be credible, researchers must justify the following assumption:
[…]
where $\Delta_{s,t}Y_{i,t}(0)=Y_{i,t}(0)-Y_{i,s}(0)$ is the before-and-after difference in untreated potential outcomes across any two periods (assuming no covariates, though this naturally extends to include them). This assumption is slightly stronger than PT, which only assumes mean independence, but is more intuitive. It demonstrates both the strength and limitation of causal panel analysis under PT: While the panel structure helps account for time-invariant unobserved confounding through before-and-after differences, the introduction of the treatment must act as a shock in so much as it is orthogonal to the evolution of untreated potential outcomes; hence, any dynamic relationships between past outcomes or covariates and treatment are ruled out. In other words, under PT, a strong causal panel design still requires some (quasi-)random element in treatment assignment. Importantly, the research design issues cannot be resolved simply by applying the novel estimators surveyed in this paper.
Inspecting raw data helps spot obvious issues.
The design phase should also include inspection of the raw data. We encourage researchers to plot the data at hand to better understand patterns of the treatment status, missingness, and outliers. Treatment status should vary both by unit and time. If the majority of variation occurs over time (across units) with little or no variation between units at any given time period (or across time within a given unit), the TWFE estimand will be likely dominated by impermissible comparisons and thus susceptible to larger biases. Moreover, HTE-robust estimators will estimate the treatment effect using very little data and thus be underpowered. Equally important is the need for researchers to understand the degree and possible origins of data missingness prior to initiating statistical analysis. If missingness does not seem to be random, or if it is too prevalent, leaving insufficient meaningful variation in the data, researchers should consider halting the analysis at this stage. Just as in the cross-sectional case, plotting the raw data can also help researchers to spot outliers and highly skewed distributions, which may require additional preprocessing. At this stage, researchers can also trim the data to make the units in comparison more similar in terms of pre-treatment or time-invariant covariates (Kahn-Lang and Lang, 2020; Sant’Anna and Zhao, 2020).
When estimation results diverge, understand why.
At the estimation stage, we recommend using at least one HTE-robust estimator alongside a benchmark TWFE model. While TWFE is often more efficient, its constant treatment effects assumption is too restrictive in many contexts. As shown in this article, TWFE does not produce systematic upward or downward bias compared to the imputation estimator, but it can severely bias causal estimates in individual cases. We recommend the imputation estimator in most settings primarily for its flexibility, though other estimators also have their advantages. The more critical question, however, is why results differ between estimators when they do. If TWFE and an HTE-robust estimator diverge, it could be due to HTE or PT violations. If HTE-robust estimators themselves diverge, it is often because the data are too sparse, leading to high variability, or the PT assumption fails differentially, causing estimators weighting control units differently to produce varying estimates. Plotting raw data and using diagnostics (such as the Goodman-Bacon decomposition) typically clarifies these issues. We also recommend keeping the benchmark TWFE model for its transparency.
For uncertainty estimates, researchers should use cluster-robust SEs when the number of clusters is large (e.g., exceeds $50$) and opt for cluster-bootstrap or cluster-jackknife procedures when the number of clusters is relatively small. The clustering level should correspond to either the level of the time-series units or the level of treatment assignment, choosing the higher of the two. For novel HTE-robust estimators, cluster-bootstrap or jackknife is generally safer than relying on various analytical SEs. In Section A.4.2 of the SM, we show that cluster-bootstrapped SEs are typically larger than analytically derived SEs for five HTE-robust estimators using data from our sample.
Conduct diagnostics to assess key assumptions and robustness of findings.
With a clear research design, researchers should critically evaluate key identification assumptions and test the robustness of findings when these assumptions are violated. Event-study plots, available for TWFE and most HTE-robust estimators, are valuable tools for assessing whether the no-anticipation and PT assumptions hold. We recommend creating an event-study plot using the chosen estimator(s), followed by both visual inspection and statistical tests to assess how plausible the PT assumption is. Importantly, the absence of statistical significance in pretrend coefficients should not be taken as evidence for the validity of PT. To avoid the conditional inference problem, we recommend performing a sensitivity analysis with robust confidence sets across different values of $\bar{M}$, regardless of the pretest results. As shown in this article, such tests require more statistical power than rejecting the null that the average treatment effect is zero. Researchers should, therefore, allocate sufficient statistical power for these diagnostic tests during the research design phase.
Panel data offer valuable opportunities for social scientists to address difficult causal questions; however, these data, especially when analyzed under PT, present specific challenges. Our findings are not meant to discourage researchers from using research designs based on PT when using panel data for causal analysis. Instead, we aim to guide them in conducting their analyses in a more transparent and credible manner. To support this effort, we have developed an open-source package that enables researchers to implement all the procedures described in this paper.
Supporting Materials: For Appendix, markdown files, and software implementation, please visit https://yiqingxu.org/research/ and look for Causal Panel Analysis under Parallel Trends: Lessons from A Large Reanalysis Study.